# Causal Classification of Pathological Misner-Type Spacetimes

### v2

@@ 3. General equations for radial geodesics
In this section, we present a complete canonical formulation of the geodesic equations for cylindrical Misner-type spacetime models, focusing on the embedded $\theta=\textrm{const}$ and $\varphi=\textrm{const}$ cylinders, as introduced in Subsection 2.2. We anticipate that the radial geodesic equations for the pseudo-Schwarzschild and pseudo-Reissner-Nordström spacetimes (to be discussed in Sections 4 and 6) will exhibit structural similarities. To unify these cases, we derive a general form of the geodesic equations, which proves to be valid for all three hyperbolically symmetric spacetimes. Notably, this general form can also be applied to Misner space, recognizing that the $2$-dimensional Misner metric, $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, can be rewritten in the canonical form
@@
To generalize the Equation 3.1 to the $4$-dimensional case for hyperbolically symmetric models, we start with the general form of a hyperbolically symmetric metric given by
@@
where $k=-1$ for timelike, $k=+1$ for spacelike, and $k=0$ for null geodesics. The time coordinate $\nu$ is cyclic, and we always have $g_{\nu r}=1$, so we obtain the constant term
@@
[…]
@@
[…]
@@
[…]
For timelike geodesics, the condition $\omega\sqrt{\xi^{2}+g_{\nu\nu}}=0\Longleftrightarrow\xi^{2}+g_{\nu\nu}=0$, with the value $-\xi^{2}=g_{\nu\nu}$, corresponds to the particle either having a turning point or asymptotically approaching a horizon.
These results provide further insights. By renaming the coordinates as $\nu=\psi$ and $r=T$, we can carry over the findings to the Misner case, which takes the same general form as the $2$-dimensional cylindrical pseudo-Schwarzschild (4.4) and pseudo-Reissner-Nordström (6.3) spacetimes in Eddington-Finkelstein coordinates:
@@ 4. Pseudo-Schwarzschild spacetime
where $S^{1}$ is the $1$-sphere, topologically equivalent to an interval whose endpoints are glued together by identification, and $H_{r}^{2}$ (when $t$, $r=\textrm{const}$) represents a $2$-dimensional surface that can be identified with one of the sheets of the two-sheeted space-like hyperboloid. [^7] The upper sheet of the hyperboloid can be globally embedded in $3$-dimensional Minkowski space. We can express the upper sheet hyperboloid using a parameterization in polar coordinates $(\theta,\phi)$. Specifically, the map +×$\mathbb{R}^{+}\times S^{1}\rightarrow H_{r}^{2}$, $(\theta,\phi)\mapsto(r{\normalcolor\cosh(\theta)},r{\normalcolor\sinh(\theta)}{\normalcolor\cos(\varphi)},r{\normalcolor\sinh(\theta)\sin(\phi)})$ describes the embedding of the upper sheet of the hyperboloid in Minkowski space. The hyperboloid together with the induced Riemannian metric $g=d\theta\otimes d\theta+\sinh^{2}(\theta)d\phi\otimes d\phi$ is called hyperbolic plane. Then the pseudo-Schwarzschild metric can be derived by performing a Wick rotation $\theta\rightarrow i\theta$ and a signature change from the famous Schwarzschild metric as follows:
@@
The pseudo-Schwarzschild spacetime is static inside the horizon and time-dependent outside of the horizon. For $r<2m$, the $t$-direction $\partial_{t}$ is timelike, and $r$ is a spatial coordinate and the $r$-direction $\partial_{r}$ is spacelike, as seen from the metric components:
@@
To eliminate the coordinate singularity at $r=2m$, we can convert the pseudo-Schwarzschild coordinates into Eddington-Finkelstein coordinates. Similar to the Schwarzschild case, we set the term $r+2m\ln\mid r-2m\mid$ equal to the tortoise coordinate $r*$. Then, we set $\nu=-(t+r*),$ which leads to the inverse relationship $t=-(\nu+r+2m\ln\mid r-2m\mid).$ Next, we take the derivative of this equation $dt=-d(\nu+r+2m\ln\mid r-2m\mid),$ and square both sides. By substituting this expression for $dt^{2}$ into the pseudo-Schwarzschild metric, we obtain the following form of the metric in terms of the tortoise coordinate
@@
The new coordinate $\nu$ is periodic, with $0\leq\nu<\beta$, similar to the previous coordinate $t$. A given point $(0,r,\theta,\phi)$ is identified with $(b,r,\theta,\phi)$, where $b$ is the periodicity of the coordinate $\nu$. These coordinates are naturally adapted to the null geodesics. Although the metric component $g_{\nu\nu}$ disappears at $r=2m$ , the coordinate transformation eliminates the singularity that existed at $r=2m$ in the original pseudo-Schwarzschild coordinates. Moreover, the null hypersurface at $r=2m$ serves as a Cauchy horizon, denoted $H$, and divides the spacetime $M$ into two regions: $M_{2}$, where $r<2m$, and $M_{1}$, where $r>2m$. We can time-orient the manifold $M$ by requiring that $-\partial_{r}$ be future-pointing. This leads to the conclusion that $\partial_{r}$ is a lightlike vector throughout, $\partial_{\nu}$ is timelike for $r<2m$, and spacelike for $r>2m$.
@@
Recalling that the pseudo-Schwarzschild spacetime is $M=M_{Z}\times_{r}H^{2}$, and its metric can be decomposed into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$, we proceed by setting $\theta=\textrm{const}$ and $\phi=\textrm{const}$. Allowing $\nu$ and $r$ to vary and take the values as previously defined, we obtain the $2$-dimensional cylindrical pseudo-Schwarzschild spacetime $M_{Z}=S^{1}\times\mathbb{R}^{+}$. This spacetime is described by the degenerate metric in Eddington-Finkelstein coordinates (see Figure 6):
@@
In two dimensions, the Christoffel symbols are identical for both the Schwarzschild and pseudo-Schwarzschild spacetimes. Therefore, the Riemann curvature tensor for the cylindrical pseudo-Schwarzschild spacetime (4.3) is given by $-R_{trtr}=-\frac{2m}{r^{3}}\eqqcolon K=\frac{1}{2}S$, where $K$ denotes the Gaussian curvature and $S$ represents the scalar curvature. Since we require $m$ to be non-negative and $r$ to be positive, the Gaussian curvature is non-positive throughout the cylinder. [^9] The Riemannian curvature tensor for the $4$-dimensional pseudo-Schwarzschild spacetime can easily computed by the prescription of Wick rotation and sign change.
@@ 4.1. Non-chronal region
In cylindrical coordinates $(\nu,r)$, the $\nu$-coordinate curves circle around the cylinder. Given the metric metric in Eddington-Finkelstein coordinates (as presented in Equation 4.4), besides $\theta,\phi$, we can also set $r$ to a constant value and consider the one-dimensional line element
@@ 4.2. Pseudo-Schwarzschild geodesics
[…]
@@ 5. Relation between pseudo-Schwarzschild and Misner spacetime
We aim to investigate Misner space and the $2$-dimensional pseudo-Schwarzschild spacetime $(M_{Z},g_{Z})$, focusing on similarities such as global causal structure, curvature, and geodesics. In addition to these, the conformal structure of these spacetimes is notably rich. A conformal transformation can sometimes convert a curved space with a non-zero Riemann tensor into a flat space with a zero Riemann tensor. However, depending on the dimension of the spacetime, a less strict concept of causal relatedness may be required. In this context, we introduce the concept of isocausality [6], which refers to a situation where the causal structure of spacetime may not be preserved globally under certain transformations, but the notion of causal relationships within specific regions remains consistent.
A glance at the causal structure of the cylindrical pseudo-Schwarzschild space and the 2-dimensional Misner space might suggest that the two associated metrics, metric (4.4) and the $2$-dimensional variant of metric (2.4), are related. This raises the natural question of whether an isometric mapping exists between these manifolds. However, it is known that any isometry of a pseudo-Riemannian manifold preserves its curvature. Since the Riemann curvature vanishes in Misner space, but does not vanish in the cylindrical pseudo-Schwarzschild space, no isometry exists between them.
@@
It is worth noting that in the special case $m=0$, the $2$-dimensional pseudo-Schwarzschild spacetime is defined on $S^{1}\times\mathbb{R}$ and is flat. However, by setting $m=0$, the chronology horizon in the pseudo-Schwarzschild space disappears, and closed timelike curves (CTCs) no longer exist. This distinction makes these spacetimes causally different, as the chronology horizon is a significant global feature that separates the chronology-violating region from the region without CTCs. In what follows, we will restrict our investigation to the case where $m>0$.
@@ 5.1. General case
For radial trajectories in hyperbolically symmetric spacetimes, the geometry is effectively $1+1$ dimensional. Therefore, for simplicity and ease of calculation, we first focus on the $2$-dimensional Misner space and the cylindrical pseudo-Schwarzschild space, as these lead to relatively straightforward conformal transformations. Afterward, we turn our attention to $4$-dimensional spacetimes, although the results from the $2$-dimensional case cannot be directly generalized. This gives rise to an alternative perspective on causal equivalence, as introduced by [6]. This approach allows for the consideration of a much larger class of causally related spacetimes compared to the classical concept of conformally related spacetimes.
@@ 5.1.1. Relation in the two-dimensional case
Since the metrics under study are $2$-dimensional, we know that both metrics in question belong to the same conformal equivalence class $[g]$ of conformally flat metrics. [^10] A metric is said to be conformally flat if it can be conformally transformed into a metric with vanishing Riemann curvature. Hence, if the metrics $g_{P},g_{M}\in[g]$ are both locally conformal to a flat metric $g$, i.e. $(g_{M}\sim g)\,\wedge\,(g_{P}\sim g)$, then by transitivity and symmetry we have $g_{P}\sim g_{M}$. As a result, we can locally transform the curved pseudo-Schwarzschild space with a non-vanishing Riemann tensor into the flat Misner space with a zero Riemann tensor by applying a conformal transformation.
Consider the $2$-dimensional flat Misner space defined on $S^{1}\times\mathbb{R}$ and equipped with the metric
@@
This version of Misner metric can be obtained from the $2$-dimensional Minkowski metric (2.1) by the initial coordinate transformation $\tilde{t}=\eta\cosh(X_{1})$ and $\tilde{x}_{1}=\eta\sinh(X_{1})$. This coordinate change results in the Misner metric $ds^{2}=-d\eta^{2}+\eta^{2}(dX_{1})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$. Another coordinate change with $\varphi=X_{1}+\ln(\eta)$ and $T=\eta^{2}$ results in the Misner metric $-dTd\varphi+Td\varphi{}^{2}$, where the domains are $-\infty<T<\infty$ and $0\leq\varphi\leq 2\pi$.
@@
$ds^{2}$ $=-(\frac{2}{\bar{r}}-1)d(\bar{\nu}m)^{2}+2d(\bar{\nu}m)d(\bar{r}m)$
$=m^{2}[-(\frac{2}{\bar{r}}-1)d\bar{\nu}^{2}+2d\bar{\nu}d\bar{r}]$
$=m^{2}(\frac{2}{\bar{r}}-1)[-d\bar{\nu}^{2}+\underset{d(r^{*})}{\underbrace{2(\frac{2}{\bar{r}}-1)^{-1}d\bar{r}}}d\bar{\nu}]$.
@@
$ds^{2}$ $=m^{2}(\frac{2}{\bar{r}}-1)[-(\frac{1}{\alpha}d\varphi)^{2}+(\frac{1}{\alpha T}dT)(\frac{1}{\alpha}d\varphi)]$
$=m^{2}(\frac{2}{\bar{r}}-1)[-\frac{1}{\alpha^{2}}d\varphi{}^{2}+\frac{1}{\alpha^{2}T}dTd\varphi]$.
By multiplication with $(-1)$ we get
$ds^{2}$ $=-m^{2}(\frac{2}{\bar{r}}-1)\frac{1}{\alpha^{2}}\frac{1}{T}[-Td\varphi{}^{2}+dTd\varphi]=\underset{\Omega(T)}{\underbrace{(\frac{2}{\bar{r}}-1)\frac{m^{2}}{T\alpha^{2}}}}[-dTd\varphi+Td\varphi{}^{2}]$
$=\Omega(T)[-dTd\varphi+Td\varphi{}^{2}]$.
@@
$\Omega(\bar{r})$ $=(\frac{2}{\bar{r}}-1)\frac{m^{2}}{\alpha^{2}(\textrm{e}^{-2\alpha\bar{r}}\cdot(2-\bar{r})^{-4\alpha})}=\frac{m^{2}}{\bar{r}\alpha^{2}}(2-\bar{r})\frac{1}{(2-\bar{r})^{-4\alpha}}\cdot\textrm{e}^{2\alpha\bar{r}}$
$=\frac{m^{2}}{\bar{r}\alpha^{2}}(2-\bar{r})(2-\bar{r})^{4\alpha}\cdot\textrm{e}^{2\alpha\bar{r}}=\frac{m^{2}\textrm{e}^{2\alpha\bar{r}}}{\bar{r}\alpha^{2}}(2-\bar{r})^{4\alpha+1}$,
@@
$\Omega(T)$ $=\frac{16m^{3}\textrm{e}^{\frac{-r}{2m}}}{r}=\frac{16m^{3}\textrm{e}^{\frac{-2m\left(W(-\frac{T}{2\textrm{e}})+1\right)}{2m}}}{2m\left(W(-\frac{T}{2\textrm{e}})+1\right)}=\frac{8m^{2}\textrm{e}^{-\left(W(-\frac{T}{2\textrm{e}})+1\right)}}{\left(W(-\frac{T}{2\textrm{e}})+1\right)}$
$=\frac{8m^{2}\textrm{e}^{-\left(W(-\frac{T}{2\textrm{e}})+1\right)}}{W(-\frac{T}{2\textrm{e}})+1}=\frac{8m^{2}}{\textrm{e}^{\left(W(-\frac{T}{2\textrm{e}})+1\right)}\cdot W(-\frac{T}{2\textrm{e}})+1}=\frac{8m^{2}}{\textrm{e}^{\left(W(-\frac{T}{2\textrm{e}})+1\right)}-\frac{T}{2}}$
@@
Conformal transformations between $2$-dimensional Lorentzian manifolds send null hypersurfaces to null hypersurfaces and therefore null geodesics to null geodesics.
@@
This means that for the $1+1$ dimensional pseudo-Schwarzschild and Misner spacetimes, we can identify conformal transformations and demonstrate that timelike geodesics are of a similar type in both spacetimes. This establishes a notable relationship between them, with the $2$-dimensional pseudo-Schwarzschild spacetime being viewed as a modification or generalization of Misner space. However, the $4$-dimensional pseudo-Schwarzschild solution is not conformally flat, as indicated by the non-zero components of the conformally invariant Weyl tensor. In fact, only one independent spin component of the Weyl tensor does not vanish, specifically the Weyl scalar $C=\frac{m}{r^{3}}$. Therefore, in this context, a less strict definition of causal relatedness is required.
5.1.2. Relation in the $4$-dimensional case
Since the Weyl curvature tensor vanishes identically in two dimensions, all $2$-dimensional Lorentz metrics belong to the same conformal class of flat metrics. Therefore, it is relatively straightforward to find conformal transformations between any two representatives of this conformal class. In dimensions greater than three, the Weyl curvature tensor is generally non-zero, and a necessary condition for a spacetime to be conformally flat is that the Weyl tensor vanishes. As a result, establishing an explicit conformal relation between $4$-dimensional spacetimes becomes significantly more challenging.

### v3

@@ 3. General equations for radial geodesics
In this section, we present a complete canonical formulation of the geodesic equations for cylindrical Misner-type spacetime models, focusing on the embedded $\theta=\textrm{const}$ and $\varphi=\textrm{const}$ cylinders, as introduced in Subsection 2.2. We anticipate that the radial geodesic equations for the pseudo-Schwarzschild and pseudo-Reissner-Nordström spacetimes (to be discussed in Sections 4 and 6) will exhibit structural similarities. To unify these cases, we derive a general form of the geodesic equations, which proves to be valid for all three hyperbolically symmetric spacetimes. Notably, this general form can also be applied to Misner space, recognizing that the 22–dimensional Misner metric, $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, can be rewritten in the canonical form
@@
To generalize the Equation 3.1 to the 44–dimensional case for hyperbolically symmetric models, we start with the general form of a hyperbolically symmetric metric given by
@@
where k=1k=-1 for timelike, $k=+1$ for spacelike, and $k=0$ for null geodesics. The time coordinate $\nu$ is cyclic, and we always have $g_{\nu r}=1$, so we obtain the constant term
@@
(3.6) $\nu(r)=\int\frac{\dot{\nu}}{\dot{r}}dr=\int\frac{g^{\nu\nu}\left(\xi-\omega\sqrt{\xi^{2}-kg_{\nu\nu}}\right)}{\omega\sqrt{\xi^{2}-kg_{\nu\nu}}}dr=\int\frac{1}{g_{\nu\nu}}\left(\frac{\omega}{\sqrt{1-\frac{kg_{\nu\nu}}{\xi^{2}}}}-1\right)dr.$
@@
(3.7) $\nu(r)=\int\frac{g^{\nu\nu}\left(\xi-\omega\sqrt{\xi^{2}+g_{\nu\nu}}\right)}{\omega\sqrt{\xi^{2}+g_{\nu\nu}}}dr=\int\frac{1}{g_{\nu\nu}}\left(\frac{\omega}{\sqrt{1+\frac{g_{\nu\nu}}{\xi^{2}}}}-1\right)dr,$
@@
(3.8) $\nu(r)=\int\frac{1}{g_{\nu\nu}}\left(\omega-1\right)dr.$
For timelike geodesics, the condition $\omega\sqrt{\xi^{2}+g_{\nu\nu}}=0\Longleftrightarrow\xi^{2}+g_{\nu\nu}=0$, with the value $\xi^{2}=g_{\nu\nu}$, corresponds to the particle either having a turning point or asymptotically approaching a horizon.
These results provide further insights. By renaming the coordinates as $\nu=\psi$ and $r=T$, we can carry over the findings to the Misner case, which takes the same general form as the 22–dimensional cylindrical pseudo-Schwarzschild (4.4) and pseudo-Reissner-Nordström (6.3) spacetimes in Eddington-Finkelstein coordinates:
@@ 4. Pseudo-Schwarzschild spacetime
where $S^{1}$ is the 11–sphere, topologically equivalent to an interval whose endpoints are glued together by identification, and $H_{r}^{2}$ (when $t$, $r=\textrm{const}$) represents a 22–dimensional surface that can be identified with one of the sheets of the two-sheeted space-like hyperboloid. [^7] The upper sheet of the hyperboloid can be globally embedded in $3$-dimensional Minkowski space. We can express the upper sheet hyperboloid using a parameterization in polar coordinates $(\theta,\phi)$. Specifically, the map $\mathbb{R}^{+}\times S^{1}\rightarrow H_{r}^{2}$, $(\theta,\phi)\mapsto(r{\normalcolor\cosh(\theta)},r{\normalcolor\sinh(\theta)}{\normalcolor\cos(\varphi)},r{\normalcolor\sinh(\theta)\sin(\phi)})$ describes the embedding of the upper sheet of the hyperboloid in Minkowski space. The hyperboloid together with the induced Riemannian metric $g=d\theta\otimes d\theta+\sinh^{2}(\theta)d\phi\otimes d\phi$ is called hyperbolic plane. Then the pseudo-Schwarzschild metric can be derived by performing a Wick rotation $\theta\rightarrow i\theta$ and a signature change from the famous Schwarzschild metric as follows:
@@
The pseudo-Schwarzschild spacetime is static inside the horizon and time-dependent outside of the horizon. For $r<2m$, the tt–direction $\partial_{t}$ is timelike, and $r$ is a spatial coordinate and the $r$-direction $\partial_{r}$ is spacelike, as seen from the metric components:
@@
To eliminate the coordinate singularity at $r=2m$, we can convert the pseudo-Schwarzschild coordinates into Eddington-Finkelstein coordinates. Similar to the Schwarzschild case, we set the term $r+2m\ln\mid r-2m\mid$ equal to the tortoise coordinate rr*. Then, we set $\nu=-(t+r*),$ which leads to the inverse relationship $t=-(\nu+r+2m\ln\mid r-2m\mid).$ Next, we take the derivative of this equation $dt=-d(\nu+r+2m\ln\mid r-2m\mid),$ and square both sides. By substituting this expression for $dt^{2}$ into the pseudo-Schwarzschild metric, we obtain the following form of the metric in terms of the tortoise coordinate
@@
The new coordinate $\nu$ is periodic, with $0\leq\nu<\beta$, similar to the previous coordinate $t$. A given point $(0,r,\theta,\phi)$ is identified with $(b,r,\theta,\phi)$, where $b$ is the periodicity of the coordinate $\nu$. These coordinates are naturally adapted to the null geodesics. Although the metric component $g_{\nu\nu}$ disappears at $r=2m$ , the coordinate transformation eliminates the singularity that existed at $r=2m$ in the original pseudo-Schwarzschild coordinates. Moreover, the null hypersurface at $r=2m$ serves as a Cauchy horizon, denoted $H$, and divides the spacetime $M$ into two regions: $M_{2}$, where $r<2m$, and $M_{1}$, where $r>2m$. We can time-orient the manifold $M$ by requiring that $\partial_{r}$ be future-pointing. This leads to the conclusion that $\partial_{r}$ is a lightlike vector throughout, $\partial_{\nu}$ is timelike for $r<2m$, and spacelike for $r>2m$.
@@
Recalling that the pseudo-Schwarzschild spacetime is $M=M_{Z}\times_{r}H^{2}$, and its metric can be decomposed into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$, we proceed by setting $\theta=\textrm{const}$ and $\phi=\textrm{const}$. Allowing $\nu$ and $r$ to vary and take the values as previously defined, we obtain the 22–dimensional cylindrical pseudo-Schwarzschild spacetime $M_{Z}=S^{1}\times\mathbb{R}^{+}$. This spacetime is described by the degenerate metric in Eddington-Finkelstein coordinates (see Figure 6):
@@
In two dimensions, the Christoffel symbols are identical for both the Schwarzschild and pseudo-Schwarzschild spacetimes. Therefore, the Riemann curvature tensor for the cylindrical pseudo-Schwarzschild spacetime (4.3) is given by $R_{trtr}=-\frac{2m}{r^{3}}\eqqcolon K=\frac{1}{2}S$, where $K$ denotes the Gaussian curvature and $S$ represents the scalar curvature. Since we require $m$ to be non-negative and $r$ to be positive, the Gaussian curvature is non-positive throughout the cylinder. [^9] The Riemannian curvature tensor for the 44–dimensional pseudo-Schwarzschild spacetime can easily computed by the prescription of Wick rotation and sign change.
@@ 4.1. Non-chronal region
In cylindrical coordinates $(\nu,r)$, the $\nu$–coordinate curves circle around the cylinder. Given the metric metric in Eddington-Finkelstein coordinates (as presented in Equation 4.4), besides $\theta,\phi$, we can also set $r$ to a constant value and consider the one-dimensional line element
@@ 4.2. Pseudo-Schwarzschild geodesics
(4.5) $\nu(r)=\int\frac{-\frac{1}{(\frac{2m}{r}-1)}\left(\xi-\omega\sqrt{\xi^{2}-(\frac{2m}{r}-1)}\right)}{\omega\sqrt{\xi^{2}-(\frac{2m}{r}-1)}}dr=\int\frac{-1}{(\frac{2m}{r}-1)}\left(\frac{\omega}{\sqrt{1-\frac{(\frac{2m}{r}-1)}{\xi^{2}}}}-1\right)dr.$
@@ 5. Relation between pseudo-Schwarzschild and Misner spacetime
We aim to investigate Misner space and the 22–dimensional pseudo-Schwarzschild spacetime $(M_{Z},g_{Z})$, focusing on similarities such as global causal structure, curvature, and geodesics. In addition to these, the conformal structure of these spacetimes is notably rich. A conformal transformation can sometimes convert a curved space with a non-zero Riemann tensor into a flat space with a zero Riemann tensor. However, depending on the dimension of the spacetime, a less strict concept of causal relatedness may be required. In this context, we introduce the concept of isocausality [6], which refers to a situation where the causal structure of spacetime may not be preserved globally under certain transformations, but the notion of causal relationships within specific regions remains consistent.
A glance at the causal structure of the cylindrical pseudo-Schwarzschild space and the 22–dimensional Misner space might suggest that the two associated metrics, metric (4.4) and the 22–dimensional variant of metric (2.3), are related. This raises the natural question of whether an isometric mapping exists between these manifolds. However, it is known that any isometry of a pseudo-Riemannian manifold preserves its curvature. Since the Riemann curvature vanishes in Misner space, but does not vanish in the cylindrical pseudo-Schwarzschild space, no isometry exists between them.
@@
It is worth noting that in the special case $m=0$, the 22–dimensional pseudo-Schwarzschild spacetime is defined on $S^{1}\times\mathbb{R}$ and is flat. However, by setting $m=0$, the chronology horizon in the pseudo-Schwarzschild space disappears, and closed timelike curves (CTCs) no longer exist. This distinction makes these spacetimes causally different, as the chronology horizon is a significant global feature that separates the chronology-violating region from the region without CTCs. In what follows, we will restrict our investigation to the case where $m>0$.
@@ 5.1. General case
For radial trajectories in hyperbolically symmetric spacetimes, the geometry is effectively $1+1$ dimensional. Therefore, for simplicity and ease of calculation, we first focus on the 22–dimensional Misner space and the cylindrical pseudo-Schwarzschild space, as these lead to relatively straightforward conformal transformations. Afterward, we turn our attention to 44–dimensional spacetimes, although the results from the 22–dimensional case cannot be directly generalized. This gives rise to an alternative perspective on causal equivalence, as introduced by [6]. This approach allows for the consideration of a much larger class of causally related spacetimes compared to the classical concept of conformally related spacetimes.
@@ 5.1.1. Relation in the two-dimensional case
Since the metrics under study are 22–dimensional, we know that both metrics in question belong to the same conformal equivalence class $[g]$ of conformally flat metrics. [^10] A metric is said to be conformally flat if it can be conformally transformed into a metric with vanishing Riemann curvature. Hence, if the metrics $g_{P},g_{M}\in[g]$ are both locally conformal to a flat metric $g$, i.e. $(g_{M}\sim g)\,\wedge\,(g_{P}\sim g)$, then by transitivity and symmetry we have $g_{P}\sim g_{M}$. As a result, we can locally transform the curved pseudo-Schwarzschild space with a non-vanishing Riemann tensor into the flat Misner space with a zero Riemann tensor by applying a conformal transformation.
Consider the 22–dimensional flat Misner space defined on $S^{1}\times\mathbb{R}$ and equipped with the metric
@@
This version of Misner metric can be obtained from the 22–dimensional Minkowski metric (2.1) by the initial coordinate transformation $\tilde{t}=\eta\cosh(X_{1})$ and $\tilde{x}_{1}=\eta\sinh(X_{1})$. This coordinate change results in the Misner metric $ds^{2}=-d\eta^{2}+\eta^{2}(dX_{1})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$. Another coordinate change with $\varphi=X_{1}+\ln(\eta)$ and $T=\eta^{2}$ results in the Misner metric $dTd\varphi+Td\varphi{}^{2}$, where the domains are $\infty<T<\infty$ and $0\leq\varphi\leq 2\pi$.
@@
$ds^{2}$ $(\frac{2}{\bar{r}}-1)d(\bar{\nu}m)^{2}+2d(\bar{\nu}m)d(\bar{r}m)$
$m^{2}[-(\frac{2}{\bar{r}}-1)d\bar{\nu}^{2}+2d\bar{\nu}d\bar{r}]$
$m^{2}(\frac{2}{\bar{r}}-1)[-d\bar{\nu}^{2}+\underset{d(r^{*})}{\underbrace{2(\frac{2}{\bar{r}}-1)^{-1}d\bar{r}}}d\bar{\nu}]$.
@@
$ds^{2}$ $m^{2}(\frac{2}{\bar{r}}-1)[-(\frac{1}{\alpha}d\varphi)^{2}+(\frac{1}{\alpha T}dT)(\frac{1}{\alpha}d\varphi)]$
$m^{2}(\frac{2}{\bar{r}}-1)[-\frac{1}{\alpha^{2}}d\varphi{}^{2}+\frac{1}{\alpha^{2}T}dTd\varphi]$.
By multiplication with (1)(-1) we get
$ds^{2}$ $m^{2}(\frac{2}{\bar{r}}-1)\frac{1}{\alpha^{2}}\frac{1}{T}[-Td\varphi{}^{2}+dTd\varphi]=\underset{\Omega(T)}{\underbrace{(\frac{2}{\bar{r}}-1)\frac{m^{2}}{T\alpha^{2}}}}[-dTd\varphi+Td\varphi{}^{2}$]
$\Omega(T)[-dTd\varphi+Td\varphi{}^{2}$].
@@
$\Omega(\bar{r})$ $(\frac{2}{\bar{r}}-1)\frac{m^{2}}{\alpha^{2}(\textrm{e}^{-2\alpha\bar{r}}\cdot(2-\bar{r})^{-4\alpha})}=\frac{m^{2}}{\bar{r}\alpha^{2}}(2-\bar{r})\frac{1}{(2-\bar{r})^{-4\alpha}}\cdot\textrm{e}^{2\alpha\bar{r}}$
$\frac{m^{2}}{\bar{r}\alpha^{2}}(2-\bar{r})(2-\bar{r})^{4\alpha}\cdot\textrm{e}^{2\alpha\bar{r}}=\frac{m^{2}\textrm{e}^{2\alpha\bar{r}}}{\bar{r}\alpha^{2}}(2-\bar{r})^{4\alpha+1}$,
@@
$\Omega(T)$ $\frac{16m^{3}\textrm{e}^{\frac{-r}{2m}}}{r}=\frac{16m^{3}\textrm{e}^{\frac{-2m\left(W(-\frac{T}{2\textrm{e}})+1\right)}{2m}}}{2m\left(W(-\frac{T}{2\textrm{e}})+1\right)}=\frac{8m^{2}\textrm{e}^{-\left(W(-\frac{T}{2\textrm{e}})+1\right)}}{\left(W(-\frac{T}{2\textrm{e}})+1\right)}$
$\frac{8m^{2}\textrm{e}^{-\left(W(-\frac{T}{2\textrm{e}})+1\right)}}{W(-\frac{T}{2\textrm{e}})+1}=\frac{8m^{2}}{\textrm{e}^{\left(W(-\frac{T}{2\textrm{e}})+1\right)}\cdot W(-\frac{T}{2\textrm{e}})+1}=\frac{8m^{2}}{\textrm{e}^{\left(W(-\frac{T}{2\textrm{e}})+1\right)}-\frac{T}{2}}$
@@
Conformal transformations between 22–dimensional Lorentzian manifolds send null hypersurfaces to null hypersurfaces and therefore null geodesics to null geodesics.
@@
This means that for the $1+1$ dimensional pseudo-Schwarzschild and Misner spacetimes, we can identify conformal transformations and demonstrate that timelike geodesics are of a similar type in both spacetimes. This establishes a notable relationship between them, with the 22–dimensional pseudo-Schwarzschild spacetime being viewed as a modification or generalization of Misner space. However, the 44–dimensional pseudo-Schwarzschild solution is not conformally flat, as indicated by the non-zero components of the conformally invariant Weyl tensor. In fact, only one independent spin component of the Weyl tensor does not vanish, specifically the Weyl scalar $C=\frac{m}{r^{3}}$. Therefore, in this context, a less strict definition of causal relatedness is required.
5.1.2. Relation in the 44–dimensional case
Since the Weyl curvature tensor vanishes identically in two dimensions, all 22–dimensional Lorentz metrics belong to the same conformal class of flat metrics. Therefore, it is relatively straightforward to find conformal transformations between any two representatives of this conformal class. In dimensions greater than three, the Weyl curvature tensor is generally non-zero, and a necessary condition for a spacetime to be conformally flat is that the Weyl tensor vanishes. As a result, establishing an explicit conformal relation between 44–dimensional spacetimes becomes significantly more challenging.